# Diffusion Timbre Transfer via Mutual Information Guided Inpainting

### v1

@@ Abstract
We study timbre transfer as an inference-time editing problem for music audio. Starting from a strong pre-trained latent diffusion model, we introduce a lightweight procedure that requires no additional training: (i) a dimension-wise noise injection that targets latent channels most informative of instrument identity, and (ii) an early-step clamping mechanism that re-imposes the input’s melodic and rhythmic structure during reverse diffusion. The method operates directly on audio latents and is compatible with text/audio conditioning (e.g., CLAP). We discuss design choices, analyze trade-offs between timbral change and structural preservation, and show that simple inference-time controls can meaningfully steer pre-trained models for style-transfer use cases. Demo available on  [^1] Audio demos: https://anon-audio-demo-25.github.io/audio_demo/
@@ 1 Introduction
Recent advances in generative modeling—most notably diffusion probabilistic models—now produce high-quality, realistic audio and offer increasingly expressive conditioning interfaces [Ho2020, Popov2021, Huang2022, Mittal2021, Huang2023, Popov2022, Choi2023]. Turning these general frameworks into a practical tool for instrument timbre transfer remains non-trivial. One prominent line trains dedicated systems for the task—often a separate model per target instrument or domain—which delivers strong quality but scales poorly and is narrowly tailored to timbre transfer [LDBridge, WaveTransfer]. A second line repurposes powerful pretrained backbones by adding trainable post-hoc control modules to expose new control channels—such as independent melodic and timbral control, but at the cost of extra training and architectural changes [hou2024stableaudiocontrol]. A third, training-free line steers sampling via guidance and related control signals, or performs per-example inference-time optimization of latent trajectories [Ho2022cfg, Hertz2022, Zhang2023, Novack2024]. While effective, these strategies typically incur per-target optimization, added compute, or model modifications—motivating methods that recover fine-grained control directly at inference within a single pretrained model and no added computational cost.
We introduce a label-aware, optimization-free method for inference-time timbre transfer. It repurposes a pretrained latent diffusion model without retraining or per-example optimization, and runs at the computational cost of standard sampling. We instantiate the approach on Diff-A-Riff [Nistal2024, Nistal2024_2], a latent diffusion system for musical accompaniment. The core challenge is that encoder latents entangle timbre (style) with structure (pitch, rhythm), and this entanglement varies by channel. We address this with a supervised mutual-information (MI) analysis against instrument labels to rank channels and derive masks. During sampling, we intervene in two ways: (i) inject noise only along timbre-aligned channels, and (ii) preserve—by clamping/restoring—structure-dominant channels across denoising steps. The result is a controlled shift in timbre toward the target while melody and rhythm remain intact.
@@ 2 RELATED WORK
Musical timbre/style transfer is often framed as domain adaptation: learn a mapping that changes instrument identity while preserving pitch and rhythm. Non-parallel cycle-consistency and related VAE–GAN variants have been adapted to musical audio [huang2019timbretron, sammut2021ttvaegan, bitton2018move]. Diffusion is increasingly used inside the same mapping paradigm—training direct spectrogram/waveform translators or connecting source→target distributions in an audio latent [comanducci2023difftransfer, WaveTransfer, LDBridge]. These models work well with a fixed, trained-on set of target domains, but typically rely on instrument labels or target-specific parameters (e.g., heads, embeddings, adapters). Moreover, achieving high fidelity and stable convergence generally requires substantial data, compute, and hyperparameter tuning.
A complementary thread exposes factorized controls so structure and timbre can be steered separately. DDSP autoencoders provide interpretable bottlenecks (F0, loudness, harmonic–noise) that support convincing swaps with little data [engel2020ddsp]. Recent work attaches lightweight control branches to large generators [wu2023musiccontrolnet, hou2024stableaudiocontrol, baker2025lilac]. These methods offer practical editing knobs while keeping a strong backbone mostly frozen, but they rely on reliable control extraction and paired control–audio during training, and bottleneck capacity/inductive bias can under-represent complex timbres and transients.
Finally, some approaches recover timbre control at inference by steering a pretrained generator’s sampling path—e.g., per-example optimization of the initial noise or time-varying inversion for stylization [novack2024ditto, li2024musicti]. These methods avoid retraining and adapt to each input, but they can be tricky to tune; optimization-based variants add computation for every example, and without explicit constraints on structure, edits can drift and distort pitch or rhythm.
@@ 3 Background
Our work builds on Diff-A-Riff (DaR) [Nistal2024_2] (Sec.3), a generative model of musical stems in the latent space of Music2Latent2 (M2L2)[Pasini2024] (Sec.3). We endow DaR with timbre transfer by performing a pre-hoc mutual-information analysis of M2L2 latents to disentangle structure from timbre-dominant dimensions, and intervene on the latter during the reverse diffusion trajectory. Although DaR natively employs EDM sampling, we also benchmark a deterministic DDIM baseline (Sec.3) for its convenient properties.
3.1 Diff-A-Riff [Nistal2024_2] is a latent diffusion model designed for musical accompaniment generation, operating in the compressed latent space of the M2L2 encoder. Following the DiT (Diffusion Transformer) architecture, it supports a wide spectrum of generation modes, from unconditional to context-guided, using various inputs like CLAP embeddings [Elizalde2023]. A key advantage over many unconditional or symbolic models is its ability to offer fine-grained structural control while maintaining stylistic flexibility. Its training on diverse datasets enables robust one-shot generation, making it an excellent foundation for structure-preserving tasks like timbre transfer.
@@
3.3 Denoising Diffusion Implicit Models (DDIM)[Song2020] generalize traditional diffusion models by allowing deterministic sampling trajectories, significantly reducing inference time without compromising generation quality. Further, DDIM enables inverting the mapping of an existing data point back to its latent representation by deterministically reversing the diffusion process. This property makes DDIM particularly suitable for tasks that require precise manipulation or preservation of structural attributes within the latent space, as in our proposed approach for timbre transfer.
@@ 4 Methodology
4.1 Dataset We use two corpora with complementary roles: (i) NSynth dataset [Engel2017] for analysis and tuning (mutual information and baseline noise calibration), and (ii) a subset of the DaR testing corpus for timbre-transfer experiments and reporting, by pairing tracks from different songs as context audio and reference audio. Each evaluation clip is 10 seconds long and contains no more than 5 seconds of silence.. We found that more than 7000 data pairs do not yield a significant variance in our evaluation metrics. We therefore adopt 7000 for parameter search.
@@ 4.2 Baseline
4.2.1 Naive Method: Partial Noise Injection (PnI) A simple method for performing timbre transfer involves naively adding Gaussian noise to the source content latents across all dimensions [Meng2022]:
$\mathbf{z}_{t_{f}}^{\text{partial}}=z_{\text{in}}+\sigma_{t_{f}}\,\epsilon,$
where $\epsilon\sim\mathcal{N}(0,I)$. The target diffusion step $t_{f}$ is determined by finding the step closest to a user-defined noise fraction $f_{\text{par}}\in(0,1]$:
@@
where $\sigma_{\max}\equiv\sigma_{T}$. Here, $\sigma_{T}$ represents the maximum noise level of the EDM schedule (at step $T$), and $f_{\text{par}}$ is determined by a user-defined noise fraction parameter. We can use these expressions to look up the discrete timestep whose noise level $\sigma_{t}$ lies closest to the desired fraction parameter $f_{\text{par}}$ (e.g. $f_{\text{par}}=0.5$ for “50 % of the schedule”). Because the EDM schedule $\sigma_{t}$ is non-linear in $t$ [Karras2022], this lookup yields a more faithful noise‐power match than a naive scaling :
$z_{\text{in}}+f_{\text{par}}*\sigma_{\max}.$
4.2.2 DDIM-based Partial Noise Injection (DDIM-PnI) We also explore a synthetic corruption with a deterministic DDIM inversion. We reuse the EDM grid $\{\sigma_{t}\}_{0..T}$ and map it to DDIM coefficients
$\alpha_{t}=\tfrac{1}{1+\sigma_{t}^{2}},\quad 1-\alpha_{t}=\tfrac{\sigma_{t}^{2}}{1+\sigma_{t}^{2}}.$
Starting from $x_{\mathrm{ctx}}$ at t=0t{=}0, we run DDIM with from noise level T and no guidance (CFG=0=0) up to $t_{f}$ (chosen via $f_{\text{par}}$), yielding $z^{\mathrm{inv}}_{t_{f}}$. Compared to the PnI ($z_{\text{in}}+\sigma_{t_{f}}\epsilon$), this variant only changes how the start state is obtained—via inversion rather than synthetic noise—while using the same schedule and step count.
4.2.3 Partial Noise-Level Selection To set the PnI and DDIM-PnI noise fraction $f_{\text{par}}$, we adopt an Ambient Diffusion Omni–style probe [Daras2025]. We train a tiny timestep-conditioned binary classifier $C_{\phi}(z_{t},t)$ to distinguish clean NSynth latents from timbre-swapped ones (swapping the timbre-related dimensions (as identified via mutual information, see Sec 4). For each candidate $f_{\text{par}}$, we corrupt latents to $t_{f}$ on the EDM grid and measure held-out accuracy. We choose the smallest $f_{\text{par}}$ at which accuracy falls to chance ($\approx 50\%$), indicating instrument ambiguity from diffusion noise alone; this yields $f_{\text{par}}^{\star}\approx 0.5$ on our dev set, and suggested a starting point for later parameter search.
@@
4.3.2 Dimension-Wise Noise Injection : Instead of drawing a fully random $z_{T}\sim\mathcal{N}(0,I)$, we construct a structured start point that already contains the target melody and rhythm but allows the model to freely re‑synthesise timbre. Given the source content waveform $x_{\mathrm{c}}$, we first obtain a latent $z^{\text{ctx}}\in\mathbb{R}^{1\times 64\times T^{\prime}}$ by running DDIM inversion for $T$ steps with the same $\sigma_{t}$ grid that will be used during generation. Compared with the pure encoder output $E(x_{\mathrm{ctx}})$ used in the baseline, $z^{\text{ctx}}$ already has the same global variance $\mathrm{Var}[z]=\sigma_{T}^{2}$ as a standard EDM start state, eliminating the per‑dimension SNR mismatch with our unique latent initialization. Next, we perform a latent partition:
@@
At the start of denoising, the initial noise state becomes:
$\mathbf{z}_{T}^{\text{tim}}$ $=\sigma_{T}\,\boldsymbol{\epsilon}\,\odot\,\mathbf{M}_{\text{timbre}}$
$\mathbf{z}_{T}^{\text{str}}$ $=z_{T}^{\text{ctx}}\,\odot\,\bigl(1-\mathbf{M}_{\text{timbre}}\bigr)$
$\mathbf{z}_{T}^{\text{dim-wise}}$ $=\mathbf{z}_{T}^{\text{tim}}+\mathbf{z}_{T}^{\text{str}}$
where $z_{T}^{\text{ctx}}$ is the $DDIM(x_{\text{in}})$ results in noisiest step T, $\boldsymbol{\epsilon}\sim\mathcal{N}(0,I)$ is freshly drawn Gaussian noise, $\sigma_{T}$ is the EDM noise levels in the starting step $T$ and $\mathbf{M}_{\text{timbre}}\in\{0,1\}^{64}$ is the binary mask that selects the channels dominant in timbre $k$. $\mathbf{z}_{T}^{\text{tim}}$ is the pure-noise initialisation of the timbre dimensions, $\mathbf{z}_{T}^{\text{str}}$ is the inverted latent of the structure dimensions, and their sum $\mathbf{z}_{T}^{\text{dim-wise}}$ is the complete latent fed to the reverse diffusion process. We run the EDM solver for steps $t=T,\dots,0$. For the first part of the trajectory
@@
The clamp-end step $t_{c}$ is chosen by $\sigma_{t_{c}}\!\approx\!f_{\text{clamp}}\,\sigma_{T}$ with a grid search, keeping a certain amount of the clean signal, allows the model to follow the overall structure during denoising, mirroring the observation that early diffusion steps fix the global structure [Ho2020]. Beyond $t_{c}$ the sampler runs unconstrained, refining fine timbre details. For the complete step-by-step denoising routine.
@@
4.5.1 Metrics: We evaluate timbre transfer quality using Fréchet Audio Distance (FAD) [Kilgour2019] and CLAP similarity [Elizalde2023]. Melodic preservation is assessed with two complementary metrics: Dynamic Pitch Distance (DPD) using the PESTO pitch-estimator [Riou2023], and note-onset F1. Note that while our qualitative demos include polyphonic examples to demonstrate robustness, objective metrics (DPD and Onset F1) were calculated strictly on monophonic segments of the test set, as PESTO is a monophonic estimator. We doubled the number of data pairs used for the parameter search to 14000 in this evaluation.

### v2

@@ Abstract
We study timbre transfer as an inference-time editing problem for music audio. Starting from a strong pre-trained latent diffusion model, we introduce a lightweight procedure that requires no additional training: (i) a dimension-wise noise injection that targets latent channels most informative of instrument identity, and (ii) an early-step clamping mechanism that re-imposes the input’s melodic and rhythmic structure during reverse diffusion. The method operates directly on audio latents and is compatible with text/audio conditioning (e.g., CLAP). We discuss design choices, analyze trade-offs between timbral change and structural preservation, and show that simple inference-time controls can meaningfully steer pre-trained models for style-transfer use cases. Demo available at [^1] Audio demos: anon-audio-demo-25.github.io/audio_demo
@@ 1 Introduction
Recent advances in generative modeling—most notably diffusion probabilistic models—now produce high-quality, realistic audio and offer increasingly expressive conditioning interfaces [12, 29, 16, 23, 15, 30, 5]. Turning these general frameworks into a practical tool for instrument timbre transfer remains non-trivial. One prominent line trains dedicated systems for the task—often a separate model per target instrument or domain—which delivers strong quality but scales poorly and is narrowly tailored to timbre transfer [21, 2]. A second line repurposes powerful pretrained backbones by adding trainable post-hoc control modules to expose new control channels—such as independent melodic and timbral control, but at the cost of extra training and architectural changes [14]. A third, training-free line steers sampling via guidance and related control signals, or performs per-example inference-time optimization of latent trajectories [13, 11, 34, 26]. While effective, these strategies typically incur per-target optimization, added compute, or model modifications—motivating methods that recover fine-grained control directly at inference within a single pretrained model and no added computational cost.
We introduce a label-aware, optimization-free method for inference-time timbre transfer. It repurposes a pretrained latent diffusion model without retraining or per-example optimization, and runs at the computational cost of standard sampling. We instantiate the approach on Diff-A-Riff [24, 25], a latent diffusion system for musical accompaniment. The core challenge is that encoder latents entangle timbre (style) with structure (pitch, rhythm), and this entanglement varies by channel. We address this with a supervised mutual-information (MI) analysis against instrument labels to rank channels and derive masks. During sampling, we intervene in two ways: (i) inject noise only along timbre-aligned channels, and (ii) preserve—by clamping/restoring—structure-dominant channels across denoising steps. The result is a controlled shift in timbre toward the target while melody and rhythm remain intact.
@@ 2 RELATED WORK
Musical timbre/style transfer is often framed as domain adaptation: learn a mapping that changes instrument identity while preserving pitch and rhythm. Non-parallel cycle-consistency and related VAE–GAN variants have been adapted to musical audio [17, 4, 3]. Diffusion is increasingly used inside the same mapping paradigm—training direct spectrogram/waveform translators or connecting source→target distributions in an audio latent [6, 2, 21]. These models work well with a fixed, trained-on set of target domains, but typically rely on instrument labels or target-specific parameters (e.g., heads, embeddings, adapters). Moreover, achieving high fidelity and stable convergence generally requires substantial data, compute, and hyperparameter tuning.
A complementary thread exposes factorized controls so structure and timbre can be steered separately. DDSP autoencoders provide interpretable bottlenecks (F0, loudness, harmonic–noise) that support convincing swaps with little data [9]. Recent work attaches lightweight control branches to large generators [33, 14, 1]. These methods offer practical editing knobs while keeping a strong backbone mostly frozen, but they rely on reliable control extraction and paired control–audio during training, and bottleneck capacity/inductive bias can under-represent complex timbres and transients.
Finally, some approaches recover timbre control at inference by steering a pretrained generator’s sampling path—e.g., per-example optimization of the initial noise or time-varying inversion for stylization [27, 20]. These methods avoid retraining and adapt to each input, but they can be tricky to tune; optimization-based variants add computation for every example, and without explicit constraints on structure, edits can drift and distort pitch or rhythm.
@@ 3 Background
Our work builds on Diff-A-Riff (DaR) [25] (Sec.3), a generative model of musical stems in the latent space of Music2Latent2 (M2L2)[28] (Sec.3). We endow DaR with timbre transfer by performing a pre-hoc mutual-information analysis of M2L2 latents to disentangle structure from timbre-dominant dimensions, and intervene on the latter during the reverse diffusion trajectory. Although DaR natively employs EDM sampling, we also benchmark a deterministic DDIM baseline (Sec.3) for its convenient properties.
3.1 Diff-A-Riff [25] is a latent diffusion model designed for musical accompaniment generation, operating in the compressed latent space of the M2L2 encoder. Following the DiT (Diffusion Transformer) architecture, it supports a wide spectrum of generation modes, from unconditional to context-guided, using various inputs like CLAP embeddings [8]. A key advantage over many unconditional or symbolic models is its ability to offer fine-grained structural control while maintaining stylistic flexibility. Its training on diverse datasets enables robust one-shot generation, making it an excellent foundation for structure-preserving tasks like timbre transfer.
@@
3.3 Denoising Diffusion Implicit Models (DDIM)[32] generalize traditional diffusion models by allowing deterministic sampling trajectories, significantly reducing inference time without compromising generation quality. Further, DDIM enables inverting the mapping of an existing data point back to its latent representation by deterministically reversing the diffusion process. This property makes DDIM particularly suitable for tasks that require precise manipulation or preservation of structural attributes within the latent space, as in our proposed approach for timbre transfer.
@@ 4 Methodology
4.1 Dataset We use two corpora with complementary roles: (i) NSynth dataset [10] for analysis and tuning (mutual information and baseline noise calibration), and (ii) a subset of the DaR testing corpus for timbre-transfer experiments and reporting, by pairing tracks from different songs as context audio and reference audio. Each evaluation clip is 10 seconds long and contains no more than 5 seconds of silence. We found that more than 7000 data pairs do not yield a significant variance in our evaluation metrics. We therefore adopt 7000 for parameter search.
@@ 4.2 Baseline
4.2.1 Naive Method: Partial Noise Injection (PnI) A simple method for performing timbre transfer involves naively adding Gaussian noise to the source content latents across all dimensions [22]: $\mathbf{z}_{t_{f}}^{partial}=z_{\text{in}}+\sigma_{t_{f}}\,\epsilon,$ where $\epsilon\sim\mathcal{N}(0,I)$. The target diffusion step $t_{f}$ is determined by finding the step closest to a user-defined noise fraction $f_{\text{par}}\in(0,1]$:
@@
where $\sigma_{\max}\equiv\sigma_{T}$. Here, $\sigma_{T}$ represents the maximum noise level of the EDM schedule (at step $T$), and $f_{\text{par}}$ is determined by a user-defined noise fraction parameter. We can use these expressions to look up the discrete timestep whose noise level $\sigma_{t}$ lies closest to the desired fraction parameter $f_{\text{par}}$ (e.g. $f_{\text{par}}=0.5$ for “50 % of the schedule”). Because the EDM schedule $\sigma_{t}$ is non-linear in $t$ [18], this lookup yields a more faithful noise‐power match than a naive scaling $z_{\text{in}}+f_{\text{par}}*\sigma_{\max}$.
4.2.2 DDIM-based Partial Noise Injection (DDIM-PnI) We also explore a synthetic corruption with a deterministic DDIM inversion. We reuse the EDM grid $\{\sigma_{t}\}_{0..T}$ and map it to DDIM coefficients $\alpha_{t}=\tfrac{1}{1+\sigma_{t}^{2}}$,  $1-\alpha_{t}=\tfrac{\sigma_{t}^{2}}{1+\sigma_{t}^{2}}$. Starting from $x_{\mathrm{ctx}}$ at t=0t{=}0, we run DDIM from noise level T and no guidance (CFG=0=0) up to $t_{f}$ (chosen via $f_{\text{par}}$), yielding $z^{\mathrm{inv}}_{t_{f}}$. Compared to the PnI ($z_{\text{in}}+\sigma_{t_{f}}\epsilon$), this variant only changes how the start state is obtained—via inversion rather than synthetic noise—while using the same schedule and step count.
4.2.3 Partial Noise-Level Selection To set the PnI and DDIM-PnI noise fraction $f_{\text{par}}$, we adopt an Ambient Diffusion Omni–style probe [7]. We train a tiny timestep-conditioned binary classifier $C_{\phi}(z_{t},t)$ to distinguish clean NSynth latents from timbre-swapped (ones swapping the timbre-related dimensions as identified via mutual information, see Sec 4). For each candidate $f_{\text{par}}$, we corrupt latents to $t_{f}$ on the EDM grid and measure held-out accuracy. We choose the smallest $f_{\text{par}}$ at which accuracy falls to chance ($\approx 50\%$), indicating instrument ambiguity from diffusion noise alone; this yields $f_{\text{par}}^{\star}\approx 0.5$ on our dev set, and suggested a starting point for later parameter search.
@@
4.3.2 Dimension-Wise Noise Injection : Instead of drawing a fully random $z_{T}\sim\mathcal{N}(0,I)$, we construct a structured start point that already contains the target melody and rhythm but allows the model to freely re‑synthesise timbre. Given the source content waveform $x_{\mathrm{ctx}}$, we first obtain a latent $z^{\text{ctx}}\in\mathbb{R}^{1\times 64\times T^{\prime}}$ by running DDIM inversion for $T$ steps with the same $\sigma_{t}$ grid that will be used during generation. Compared with the pure encoder output $E(x_{\mathrm{ctx}})$ used in the baseline, $z^{\text{ctx}}$ already has the same global variance $\mathrm{Var}[z]=\sigma_{T}^{2}$ as a standard EDM start state, eliminating the per‑dimension SNR mismatch with our unique latent initialization. Next, we perform a latent partition:
@@
At the start of denoising, the initial noise state becomes: $\mathbf{z}_{T}^{\text{tim}}=\sigma_{T}\,\boldsymbol{\epsilon}\,\odot\,\mathbf{M}_{\text{timbre}};\quad\mathbf{z}_{T}^{\text{str}}=z_{T}^{\text{ctx}}\,\odot\,\bigl(1-\mathbf{M}_{\text{timbre}}\bigr);\quad\mathbf{z}_{T}^{\text{dim-wise}}=\mathbf{z}_{T}^{\text{tim}}+\mathbf{z}_{T}^{\text{str}},$ where $z_{T}^{\text{ctx}}$ is the $DDIM(x_{\text{ctx}})$ results in noisiest step T, $\boldsymbol{\epsilon}\sim\mathcal{N}(0,I)$ is freshly drawn Gaussian noise, $\sigma_{T}$ is the EDM noise levels in the starting step $T$ and $\mathbf{M}_{\text{timbre}}\in\{0,1\}^{64}$ is the binary mask that selects the channels dominant in timbre $k$. $\mathbf{z}_{T}^{\text{tim}}$ is the pure-noise initialisation of the timbre dimensions, $\mathbf{z}_{T}^{\text{str}}$ is the inverted latent of the structure dimensions, and their sum $\mathbf{z}_{T}^{\text{dim-wise}}$ is the complete latent fed to the reverse diffusion process. We run the EDM solver for steps $t=T,\dots,0$. For the first part of the trajectory ($t\geq t_{c}$), we clamp the structure dimensions:
@@
The clamp-end step $t_{c}$ is chosen by $\sigma_{t_{c}}\!\approx\!f_{\text{clamp}}\,\sigma_{T}$ with a grid search, keeping a certain amount of the clean signal, allows the model to follow the overall structure during denoising, mirroring the observation that early diffusion steps fix the global structure [12]. Beyond $t_{c}$ the sampler runs unconstrained, refining fine timbre details.
@@
4.5.1 Metrics: We evaluate timbre transfer quality using Fréchet Audio Distance (FAD) [19] and CLAP similarity [8]. Melodic preservation is assessed with two complementary metrics: Dynamic Pitch Distance (DPD) using the PESTO pitch-estimator [31], and note-onset F1. Note that while our qualitative demos include polyphonic examples to demonstrate robustness, objective metrics (DPD and Onset F1) were calculated strictly on monophonic segments of the test set, as PESTO is a monophonic estimator. We doubled the number of data pairs used for the parameter search to 14000 in this evaluation.